Find (f+g)(x) for the following two functions:
f(x)=2x+1
g(x)=5x-3

Answer by geetha_rama(94)

(f+g)x = f(x) + g(x) (Addition of two real function)
(f+g)x = 2x +1+5x -3
= 3x - 2